OpenStudy (anonymous):

The perimeter of a rectangle is 80inches. If the length of the rectangle is five inches more than four times the width, find the dimensions.

OpenStudy (ria23):

uhm, I'm not totally sure I have it, but I'm pretty sure yhu would set it up the way it's worded. 80=5+4w or x. whatever variable yhu decide to use doesn't matter.

OpenStudy (ria23):

Can someone correct me if I'm wrong?

OpenStudy (anonymous):

the perimeter of a rectangle formula is P= 2(l+w)

OpenStudy (johnweldon1993):

P = 2w + 2l L = 4W + 5 looks like a system of equations

OpenStudy (johnweldon1993):

so you would just substitute what we know 'L' equals..and put that into the first equation for every 'l' we see P = 2w + 2l Perimeter = 80 btw...so 80 = 2w + 2l will now become 80 = 2w + 2(4w + 5) Distribute that 2 80 = 2w + 8w + 10 Subtract 10 from both sides of this equation 70 = 2w + 8w combine like terms... 70 = 10w Divide both sides by 10 w = 7 If we know w = 7 ...we can solve for l by plugging that into either equation we have 80 = 2w + 2l 80 = 2(7) + 2l 80 = 14 + 2l 64 = 2l l = 32 Now is that correct? We knew that L = 4W + 5 If that is true lets plug in W and L 32 = 4(7) + 5 32 = 28 + 5 32 = 32 so yes it works out... w = 7 and L = 32

OpenStudy (ria23):

Dang.

OpenStudy (johnweldon1993):

Lol it just looks like a bit of work...but it's not too bad :)

OpenStudy (ria23):

Haha, system of equations is my worst area. ;n ;

OpenStudy (johnweldon1993):

Well if you ever need help, message me @Ria23 :) I'll explain it out step by step like I did here :D

OpenStudy (ria23):

I have a question that I need answered, it's median-median line, do yhu think yhu might be able to help me?